OpenStudy (anonymous):

which is an equation of a circle with center (2,-1) that passes through the point (3,4)

OpenStudy (amistre64):

do you recall the general equation for a circle?

OpenStudy (anonymous):

what do you mean?

OpenStudy (amistre64):

well, in order to construct the equation of a circle, its useful to remember the general equation of a circle from your material

OpenStudy (amistre64):

it tends to be the distance formula of the 2 points, and then you generalize the center

OpenStudy (anonymous):

im not exactly sure what you mean. that is all the question says..

OpenStudy (amistre64):

the question does not exist in a vaccuum ... it is related to the material that you are spose to be studying

OpenStudy (amistre64):

the material that you are spose to be studying would have presented you with the general equation of a circle; to which you would then be applying that general equation with the specifics stated in the problem

OpenStudy (anonymous):

these are the options: a) (x + 2)2 + (y - 1)2 = 26 b (x - 2)2 + (y + 1)2 = 13 c) (x + 2)2 + (y - 1)2 = 13 d) (x - 2)2 + (y + 1)2 = 26

OpenStudy (amistre64):

those are good options :) .... one of them is the proper equation when the general setup is filled in ... the general setup of a circle with radius (r) and center (h,k) is defined as:\[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

it doesnt give me the radius though. can you please help me out im still extremely confused....

OpenStudy (amistre64):

the radius is the distance from the center to a given point; which can be formulated as:\[r^2=(x_o-x_1)^2+(y_o-y_1)^2\]

OpenStudy (anonymous):

is it b?

OpenStudy (amistre64):

close .. but no. -2 and +1 are fine, but im getting 26 instead of 13

OpenStudy (anonymous):

its d right :)

OpenStudy (amistre64):

i like d yes :)

OpenStudy (anonymous):

Thank you sooooooo much!

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

thanks :)))